OpenStudy (anonymous):

Hey, can anyone help me with this calculus problem Differentiate the function. y = e^(x + 2) + 4 y' = ?

OpenStudy (anonymous):

\[\frac{ d(e ^{u}) }{ dx } = e ^{u} \times \frac{ du }{ dx }\]Just determine "u".

OpenStudy (anonymous):

\[y = e ^{x} e ^{2} + 4\] Now differentiate.

OpenStudy (anonymous):

e^(x+2)

OpenStudy (anonymous):

use the chain rule

OpenStudy (anonymous):

Do you see that u = x + 2 ? So, make the substitution and simplify.

OpenStudy (anonymous):

Thanks for the replys everyone. I'm struggling with calculus and not totally sure how to do many problems. So boolan_abhi. This isn't about the power rule?

OpenStudy (anonymous):

For functions where you have "e" to a given power, you just use the formula from my first post.

OpenStudy (anonymous):

Power rule is used when it is of the form x power something. Here we have e power something. So follow the first post of tcarrol010.

OpenStudy (anonymous):

if u=g(x) is differentiable at x and y=f(u) is differentiable at u, then y is differentiable at x i.e. dy/dx=(dy/du)*(du/dx)

OpenStudy (kirbykirby):

It's actually using the chain rule, not the power rule, which is used for x power something, as @koushik_ksv said

OpenStudy (anonymous):

One other thing, the second term is a constant , "4", so you can drop that right off if you want to because the derivative of a constant is "0".

OpenStudy (anonymous):

okay. So here is what I have. I'm not sure if this is right though. e^(x+2) * (x+2)/(dx) ? What ix dx

OpenStudy (anonymous):

d(x+2)/(dx) not (x+2)/(dx) d(x+2)/(dx) = 1

OpenStudy (anonymous):

In words: the derivative with respect to "x" of "x + 2" is 1.

OpenStudy (anonymous):

I'm extremely confused with all these terms

OpenStudy (anonymous):

When taking the derivative of "x + 2", you are taking the derivative of "x" plus the derivative of "2". The derivative of "2" is "0".

OpenStudy (anonymous):

okay. i kinda get that. where does e come into this?

OpenStudy (anonymous):

Have you heard of "e" before? It's the base for natural logs.

OpenStudy (anonymous):

i heard of it before. but i don't understand the concept of it besides that it's a number

OpenStudy (anonymous):

You are taking the derivative of an exponential function. It doesn't always have to be "e", but when it is the base, derivatives are greatly simplified for exponential functions.

OpenStudy (anonymous):

derivative of e^x is e^x . WE can proof it usingfirst principle. Let, d(a^x)/dx=log a put a=e d(e^x)/dx=e^x*(ln e)=e^x

OpenStudy (anonymous):

Okay, since derivative of e^x is e^x (using chain rule?) and constant is 0. Then the derivative would be e^x?

OpenStudy (anonymous):

You are looking for the derivative of e^(x + 2) not e^x You don't have to re-create the wheel by expanding e^(x+ 2) into (e^x)(e^2). That's wasted time. The formula has already been developed and it is much quicker just to use the formula from my first post which people who use calculus eventually commit to memory.

OpenStudy (anonymous):

yup i go wid t carrol but its always wise to use first principles ..

OpenStudy (anonymous):

In words, when you have "e" to some power, the derivative is that given "e" to some power TIMES the derivative of the exponent. That's all there is to it.

OpenStudy (anonymous):

So, what's the exponent?

OpenStudy (anonymous):

(x+2)?

OpenStudy (anonymous):

Yes, and what is the derivative of that exponent. Just the exponent. In other words, what is the derivative of: x + 2 ?

OpenStudy (anonymous):

If you don't know the derivative of: x + 2 then I have to refer you to studying the basics of derivatives before trying to do the problems. You will have to know that first.